OpenStudy (anonymous):

evaluate -5x^2-3x for x=-2

Miracrown (miracrown):

Would you like to try this one first?

jimthompson5910 (jim_thompson5910):

I'll show you how to evaluate -5x^2-3x for x = -7 -5x^2-3x -5(-7)^2-3(-7) ... replace EVERY x with -7 -5(49)-3(-7) -245+21 -224 So that means -5x^2-3x = -224 when x = -7 The value -224 is NOT the answer because it's just a similar example to show you how this all works.